OpenStudy (anonymous):

Which is the equation of the line that passes through (10, 6) and (5, 3)?

OpenStudy (rabya!):

Are you aware of the formula ?

OpenStudy (anonymous):

Well i found the slope, -3/15

OpenStudy (anonymous):

its it y-3=-3/15(x-5)

OpenStudy (rabya!):

yes now simplify it.

OpenStudy (anonymous):

my answer isnt listed in the choices

OpenStudy (rabya!):

what are the choices ?

OpenStudy (rabya!):

did you copy the question right ?

OpenStudy (anonymous):

yes

OpenStudy (rabya!):

how did you find out slope ?

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (rabya!):

m= y2-y1 / x2-x1

OpenStudy (anonymous):

i did that

OpenStudy (rabya!):

how ?

OpenStudy (anonymous):

3-6/5--10

OpenStudy (anonymous):

its -10, not 10 sorry

OpenStudy (anonymous):

slope= -3/15

OpenStudy (rabya!):

m = 3-6/ 5-10 m= -3/-5 m=3/5 this should be the slope.

OpenStudy (anonymous):

Its -10

OpenStudy (anonymous):

not 10, i copied the 10 wrong

OpenStudy (anonymous):

the only option for 3/5 is negative so that isnt right

OpenStudy (rabya!):

ok.. so y-y1 = -3/15 (x-x1) Is that right ?

OpenStudy (anonymous):

yes

OpenStudy (rabya!):

so it should be y-6 = -3/15 (x+10)

OpenStudy (anonymous):

yes

OpenStudy (rabya!):

or y-6 = -1/5 ( x+10)

OpenStudy (rabya!):

right ?

OpenStudy (anonymous):

how did you get 1/5

OpenStudy (rabya!):

divide 3x1 =3 and 3x5 = 15 so simplify it 3/15 = 1/5

OpenStudy (anonymous):

Oh.. haha ok thanks

OpenStudy (rabya!):

I'm sorry but i can't figure it out